# Learning in Strategic Queuing Systems with Small Buffers

### v1

@@ 3 Analysis
By condition 2 the total number of packets served is at least $\sum_{j}(\frac{1}{2}\mu_{j}-\delta)T$. By condition 1 the total number of arriving packets is at most $\sum_{i}(\lambda_{i}+\delta)T$. Combining these two bounds establishes the lemma. ∎
Lemma 5.
Suppose there is a server whose buffer is empty more than half of the $T$ iterations, and we are in the goodδ case. Consider a queue $i$ and assume that $\delta<\frac{1}{2}(\frac{1}{2}-\lambda_{i})$, and $\delta<\lambda_{i}$. Then, the number of packets at this queue will decrease, unless the number $N^{\prime}_{i}$ left at the end the period is at most $N^{\prime}_{i}\leq(\lambda^{2}_{i}+3\delta)T$. If at the start there are at least $T$ packets in the queue, then the number decreases by at least $\left(\frac{1}{2}-(\lambda_{i}+2\delta)\right)T$.
Proof.
Consider a queue $i$. Suppose that the queue starts with $N_{i}$ packets and ends the period with $N^{\prime}_{i}$. If the server was sending a packet in every step, then by condition 3 it cleared at least $(\frac{1}{2}-\delta)T$ packets, and by condition 1 at most $(\lambda_{i}+\delta)T$ arrived, decreasing the total by at least $(\frac{1}{2}-(\lambda_{i}+2\delta))T$. This proves the second part of the claim, as $T$ packets (or more) at the start will guarantee sending a packet in every step.
Now suppose that the queue has packets in the last $\hat{T}$ steps but was empty in the previous step. Let $S_{i}$ denote the number of packets it successfully cleared in this $\hat{T}$-step period. Assume that some $(T-\hat{T})\lambda_{i}+X_{i}$ packets arrived during the first $(T-\hat{T})$ steps with $X_{i}\leq\delta T$ by Condition 1. Now by the no-regret Condition 3 we get that
$N_{i}+\lambda_{i}(T-\hat{T})+X_{i}+S_{i}\geq\frac{1}{2}(N_{i}+\lambda_{i}(T-\hat{T})+X_{i}+\hat{T})-\delta T,$
which implies (using that $X_{i}\leq\delta T$)
$N_{i}\geq\hat{T}-\lambda_{i}(T-\hat{T})-2S_{i}-3\delta T.$
By condition 1 the final number of packets in the queue at the end of the period is at most $N^{\prime}_{i}\leq\lambda_{i}\hat{T}+\delta T-S_{i}$.
Putting these two inequalities together we get that
$N_{i}-N^{\prime}_{i}$$\geq$ $\hat{T}-\lambda_{i}(T-\hat{T})-2S_{i}-3\delta T$
$-(\lambda_{i}\hat{T}+\delta T-S_{i})$
$\geq$ $\hat{T}-\lambda_{i}T-4\delta T-S_{i}.$
This shows that the number of packets at queue $i$ decreases, unless $\hat{T}\leq(\lambda_{i}+4\delta)T+S_{i}$. However, in this case we get that the number of packets at the end of the period is at most $N_{i}^{\prime}\leq\lambda_{i}\hat{T}+\delta T-S_{i}\leq(\lambda^{2}_{i}+4\lambda_{i}\delta)+\delta)T-(1-\lambda_{i})S_{i}\leq(\lambda^{2}_{i}+3\delta)T$, as $\lambda_{i}<1/2$, proving the claim. ∎
Now we are ready to prove Theorem 1.
Proof.
@@
(Theorem 1) Recall the potential function $\Phi=\sum_{i}(N_{i}-(\lambda^{2}_{i}+3\delta)T)^{+}$. Clearly, $\Phi$ has bounded moments, as the maximum possible increase in the potential function over a period of length $T$ is when a packet arrives at every single queue during each step and no packets are accepted at any of the servers. That is a total increase of $nT$. Similarly, the maximum possible decrease in the potential function over a period of length $T$ is when no packets arrive, and each queue succeeds in sending a packet to a server at each step. That is a total decrease of $nT$.
@@
Next, we want to prove that when $\Phi>Tn$, the potential is expected to decrease. Recall that when goodδ fails to hold the potential can increase by as much as $nT$. By Lemmas 2 and 3 this can contribute at most $\eta(m+2n)nT$ to the expectation.
@@
To analyze the case when goodδ holds we consider two cases separately. If all servers have packets to try to serve (in their buffer, or just arriving) at least half of the time, then by Lemma 4 the total number of packets decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$. Some of this decrease may not affect the potential function, since some of the cleared packets can come from queues that contribute zero to the potential. However, this can only affect $\sum_{i}(\lambda^{2}_{i}+3\delta)T$ packets in queue $i$, and so the potential is decreasing by at least
$\Big{(}\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}(\lambda_{i}+\lambda_{i}^{2})-(4n+m)\delta\Big{)}T.$ (1)
@@
By the assumption that $\lambda_{i}<\frac{1}{2}$ and $\sum_{j}\mu_{j}>3\sum_{i}\lambda_{i}$, this is an $\Omega(T)$ decrease, assuming we choose a small enough $\delta$.
@@
Now consider the case where a server is open (i.e., its buffer is empty, and no new packet arrived) half the time. If $\Phi>nT$ then some queue $i$ will have at least $T$ packets. In this case, by Lemma 5 this queue will have its packet count decrease by at least $(\frac{1}{2}-\lambda_{i}-2\delta)T$, contributing this decrease to the potential function. Lemma 5 also shows that all other queues will have their contribution to the potential decrease, unless it is zero at the end of the $T$-length period. This shows a total decrease in the potential of at least $(\frac{1}{2}-\lambda_{i}-2\delta)T=\Omega(T)$, again assuming $\delta$ is small enough.
@@
Putting together the expected increase when goodδ fails and expected decrease in the two cases when goodδ holds (and using that $\lambda_{i}\leq\frac{1}{2}$), we get that the potential function is expected to decrease after a $T$-length period by at least
$T\min\big{(}\frac{1}{2}-\lambda_{i}-2\delta,\frac{1}{2}\sum_{j}\mu_{j}-\frac{3}{2}\sum_{i}\lambda_{i}-(4n+m)\delta\big{)}$
$-T\eta(m+2n)n.$
@@
To guarantee the required expected decrease of the potential, we choose $\delta$ small enough so both terms in the $\min$ are at least a constant (so in the case goodδ happens the potential decreases by $\Omega(T)$) and then choose $\eta$ small enough to make the total positive. ∎
Remark 1.
Note that the proof actually shows a slightly stronger statement dependent on the $\lambda_{i}$ values. Using equation (1), it is enough to require $\sum_{i}(\lambda_{i}+\lambda_{i}^{2})<\frac{1}{2}\sum_{j}\mu_{j}$ in place of $\sum_{i}\lambda_{i}<\frac{1}{3}\sum\mu_{j}$. This improves the bound when the arrival rates are small, approaching a factor of 2 in this case.
@@
While we do not know if the factor of 3 in the main theorem is a tight worst-case bound even when arrival rates $\lambda_{i}$ are close to $\frac{1}{2}$ (in which case $\sum_{i}(\lambda_{i}+\lambda_{i}^{2})=\sum_{i}\lambda_{i}(1+\lambda_{i})=\frac{3}{2}\sum_{i}\lambda_{i}$ yields this factor), the following example shows that we need at least a factor of 2 to ensure stability.
Theorem 3.
@@
[Lower Bound Example] Consider a queue with $1/2$ arrival rate, facing $k$ servers, each with only capacity $1/n$. Assume that the server is choosing a uniform random server at each step. This sending plan will clearly satisfy the no-regret condition of Condition 3. We claim that we need at least $n$ servers to make the solution stable proving a lower bound of 2 needed for the needed increase in server capacity.
@@
$\sum_{i=1}^{\infty}1/k\cdot(1-1/k)^{i-1}(1-(1-1/n)^{i})$
$=$ $\frac{1}{k}\sum_{i=1}(1-\frac{1}{k})^{i-1}-\frac{n-1}{nk}\sum_{i=1}^{\infty}((1-\frac{1}{k})(1-\frac{1}{n}))^{i-1}$
$=$ $1-\frac{n-1}{nk}\frac{1}{1-(1-1/k)(1-1/n)}=1-\frac{n-1}{n+k-1}$
To make this stable we need this value to be above $1/2$, so we need $k>n-1$. This results in a total capacity of $\frac{k}{n}>\frac{n-1}{n}$, essentially double the arrival rate of the queue. ∎
@@ 4 Experiments
In the preceding section, we derived theoretical stability conditions and worst-case bounds on the buildup of queue lengths. Specifically, our proof showed that in a system with singleton buffers, whenever a queue grows large, given that a condition of the total service capacity is satisfied, a no-regret property of the server-selection process guarantees that it will shrink at a linear rate, keeping the total lengths of queues bounded. In this section, we complement these theoretical guarantees with computational experiments to observe the typical behavior of these systems beyond the worst case, and explore the empirical impact of buffers on the dynamics.
We conduct simulations using the EXP3 algorithm Auer et al. (2002b) (see also Slivkins (2019)), implemented in Python. For the simulation parameters, we use $\gamma$ for exploration rate and $T$ for the time horizon of the simulation.
@@ 4.1 Empirical Queue Buildup
(a) Probability that a randomized system with time horizon $T$ has a queue with buildup greater than $\sqrt{T}$ as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$. Shaded regions show the 95% binomial-proportion confidence interval for the probability estimation.
@@
(a) Dynamics in a system with $2$ queues with $\lambda_{1}=1/3$, $\lambda_{2}=1/6$ and 3 servers with $\mu_{1}=2/3$, $\mu_{2}=2/9$, $\mu_{3}=1/9$.
(b) Probability of sending to server 1 in a system with two queues with $\lambda_{i}=1/4$, and two servers with $\mu_{i}=2/3$, compared to the Nash equilibria. Dashed lines are the time average.
@@
We start by examining the relationship between the system’s arrival rate and its service capacity. The service capacity must strictly exceed the arrival rate to prevent queue buildup. This observation is seen also in typical experimental scenarios.
Figure 2 illustrates the total empirical buildup, which is the sum of all packets left in queues after $T$ iterations, normalized by $n\cdot T$ (vertical axis) as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (horizontal axis). Figure 2(a) focuses on a symmetric system with three servers such that $\mu_{1}=0.8$ and $\mu_{2}=\mu_{3}=0.2$ and three queues with equal arrival rates $\lambda_{i}$, which range between $0.02$ and $0.4$. The simulations run for $T=50$,$000$ steps with $\gamma=1/\sqrt{T}$. The solid line represents the average buildup over $200$ independent simulations for each value of $\lambda_{i}$, and the shaded region indicates the range between minimum and maximum buildup values observed in the simulations. A clear transition emerges: when $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}>0.9$, the buildup becomes proportional to $T$. Crucially, this transition occurs at a point above $0.5$ and below $1$ (where buildup is inevitable).
Figure 2(b) presents the same analysis for an ensemble of randomized systems with $5$ queues and $6$ servers. To generate such instances, we proceed as follows: generate $200$ base parameters by selecting 5 queue arrival rates, $\lambda_{i}$, and 6 server capacity rates, $\mu_{j}$, uniformly at random in $(0,1)$. For each value of the ratio $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (with $r$ in $(0,1]$), rescale the arrival rates such that the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}=r$ is satisfied. This process is repeated for each of the $200$ sets of base parameters. The simulations are run for $T=50$,$000$ and $\gamma=1/\sqrt{T}$. In Figure 2(b), the solid line represents the average buildup over the $200$ simulation for each $r$ value; the shaded region indicates the range between the $2.5$ and $97.5$ percentiles of buildup values. The results confirm that, even for asymmetric and randomized systems, the transition point remains above $0.5$ and below $1$. In both Figure 2(a) and Figure 2(b) the vertical lines indicate where $r=\frac{1}{3}$ and $r=\frac{1}{2}$. The empirical buildups in these simulations show that while the worst case bound is $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, systems with learning typically do better than the worst case at clearing packets.
@@ 4.2 The Impact of Buffers
We now compare systems with vs. without buffers. The methodology follows the one described in the previous subsection. Figure 3(a) illustrates the effect of adding buffers to a system in the random-instance scenario, with $n=5$ queues and $m=6$ servers as described earlier. It shows the probability that some queue in the system has at time $T$ a buildup greater than $\sqrt{T}$ as a function of the arrival-to-capacity ratio for systems with and without buffers. The solid lines show the observed frequency in $300$ independent simulations for each value of $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ in $(0,1]$; the shaded region shows the $95\%$ confidence interval in estimating these probabilities. We observe that systems without buffers require a higher capacity relative to the arrival rate to prevent buildup.
Figure 3(b) shows the distributions of packet clearing rates for two identical systems, with and without buffers, over $1$,$000$ simulations with $T=10$,$000$ and $\gamma=1/\sqrt{T}$. There are $4$ queues with $\lambda_{1}=0.6,\lambda_{2}=\lambda_{3}=\lambda_{4}=0.3$ and $5$ servers with $\mu_{1}=0.8,\mu_{2}=\mu_{3}=0.4,\mu_{4}=\mu_{5}=0.2$. The clear separation between the distributions demonstrates that buffers increase efficiency. In the system without buffers, queues preferred the servers with higher capacity, which led to more competition and a lower clearing rate since the three high-capacity servers could not process all packets. [^4] The total service rate of the three highest-capacity servers is exactly the same as the total arrival rate, hence not enough for stability. In the system with buffers, queues do learn to extract good service rates also from the lower capacity servers, which increases the overall clearing rate of the system.

### v2

@@ 3 Analysis
(Theorem 1) Recall the potential function $\Phi=\sum_{i}(N_{i}-(\frac{1}{2}\lambda_{i}+2\delta)T)^{+}$. Clearly, $\Phi$ has bounded moments, as the maximum possible increase in the potential function over a period of length $T$ is when a packet arrives at every single queue during each step and no packets are accepted at any of the servers. That is a total increase of $nT$. Similarly, the maximum possible decrease in the potential function over a period of length $T$ is when no packets arrive, and each queue succeeds in sending a packet to a server at each step. That is a total decrease of $nT$.
@@
Next, we want to prove that when $\Phi>Tn$, the potential is expected to decrease. Recall that when goodδ fails to hold the potential can increase by as much as $nT$. By Lemmas 4 and 5, this can contribute at most $\eta(m+2n)nT$ to the expectation.
@@
To analyze the case when goodδ holds, we consider two cases separately. If all servers have packets to try to serve (in their buffer, or just arriving) at least half of the time, then by Lemma 2 the total number of packets decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$. Some of this decrease may not affect the potential function, since some of the cleared packets can come from queues that contribute zero to the potential. However, this can only affect $\sum_{i}(\frac{1}{2}\lambda_{i}+2\delta)T$ packets, and at most $(\frac{1}{2}\lambda_{i}+2\delta)T$ packets in queue $i$, and so the potential is decreasing by at least
$\Big{(}\frac{1}{2}\sum_{j}\mu_{j}-\frac{3}{2}\sum_{i}\lambda_{i}-(3n+m)\delta\Big{)}T.$
@@
By the assumption that $\sum_{j}\mu_{j}>3\sum_{i}\lambda_{i}$, this is an $\Omega(T)$ decrease, if we choose a small enough $\delta$.
@@
Now consider the case where a server is open (i.e., its buffer is empty, and no new packet arrived) half the time. If $\Phi>nT$ then some queue $i$ will have at least $T$ packets. In this case, by Lemma 3, this queue will have its packet count decrease by at least $(\frac{1}{2}-\lambda_{i}-2\delta)T$, contributing this decrease to the potential function. Lemma 3 also shows that all other queues will have their contribution to the potential decrease, unless it is zero at the end of the $T$-length period. This shows a total decrease in the potential of at least $(\frac{1}{2}-\lambda_{i}-2\delta)T=\Omega(T)$, again assuming $\delta$ is small enough.
@@
Putting together the expected increase when goodδ fails to hold and expected decrease in the two cases when goodδ holds (and using that $\lambda_{i}<\frac{1}{2}$), we get that the potential function is expected to decrease after a $T$-length period by at least
$T\min(\frac{1}{2}-\lambda_{i}-2\delta,\ \frac{1}{2}\sum_{j}\mu_{j}-\frac{3}{2}\sum_{i}\lambda_{i}-(3n+m)\delta)-T\eta(m+2n)n.$
@@
To guarantee the required expected decrease of the potential, we choose $\delta$ small enough so that both terms in the $\min$ are at least a positive constant (so in the case goodδ happens the potential decreases by $\Omega(T)$) and then choose $\eta$ small enough to make the total positive. ∎
@@
While we do not know if the factor of 3 in the main theorem is a tight worst-case bound, the following example shows that we need at least a factor of 2 to ensure stability.
Theorem 3 (Lower Bound).
@@
Consider a queue with $1/2$ arrival rate, facing $k$ servers, each with only capacity $1/n$. Assume that the server chooses a uniform random server at each step. This sending plan will clearly satisfy the no-regret Condition 3. We claim that we need at least $n$ servers to make the solution stable proving a lower bound of 2 needed for the needed increase in server capacity.
@@
$\sum_{i=1}^{\infty}1/k\cdot(1-1/k)^{i-1}(1-(1-1/n)^{i})=\frac{1}{k}\sum_{i=1}(1-\frac{1}{k})^{i-1}-\frac{n-1}{nk}\sum_{i=1}^{\infty}\big{(}(1-\frac{1}{k})(1-\frac{1}{n})\big{)}^{i-1}$
$=1-\frac{n-1}{nk}\frac{1}{1-(1-1/k)(1-1/n)}=1-\frac{n-1}{n+k-1}.$
To make this stable, we need this value to be above $1/2$, so we need $k>n-1$. This results in a total capacity of $\frac{k}{n}>\frac{n-1}{n}$, essentially double the arrival rate of the queue. ∎
@@ 4 Experiments
In the preceding section, we derived theoretical stability conditions and worst-case bounds on the buildup of queue lengths. Specifically, our proof showed that in a system with singleton buffers, whenever a queue grows large, given that a condition of the total service capacity is satisfied, a no-regret property of the agents selecting servers guarantees that it will shrink at a linear rate, keeping the total lengths of queues bounded. In this section, we complement these theoretical guarantees with computational experiments to observe the typical behavior of these systems beyond the worst case, and explore the empirical impact of buffers on the game dynamics.
We conduct simulations using the EXP3 algorithm Auer et al. (2002b) (see also Slivkins (2019)), implemented in Python. For the simulation parameters, we denote $\gamma$ for exploration rate and $T$ for the time horizon of the simulation.
@@ 4.1 Empirical Queue Buildup
We start by studying the relationship between the system’s arrival rate and its service capacity. The service capacity must strictly exceed the arrival rate to prevent queue buildup. This observation is seen also in typical experimental scenarios.
Figure 2 illustrates the total empirical buildup, which is the sum of all packets left in queues after $T$ iterations, normalized by $n\cdot T$ (vertical axis) as a function of the arrival-to-capacity ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (horizontal axis). Figure 2(a) focuses on a symmetric system with three servers such that $\mu_{1}=0.8$ and $\mu_{2}=\mu_{3}=0.2$ and three queues with equal arrival rates $\lambda_{i}$, which range between $0.02$ and $0.4$. The simulations run for $T=50$,$000$ steps with $\gamma=1/\sqrt{T}$. The solid line represents the average buildup over $200$ independent simulations for each value of $\lambda_{i}$, and the shaded region indicates the range between minimum and maximum buildup values observed in the simulations. A clear transition emerges: when $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}>0.9$, the buildup becomes proportional to $T$. Crucially, this transition occurs at a point above $0.5$ and below $1$ (where buildup becomes inevitable).
Figure 2(b) presents the same analysis for an ensemble of randomized systems with $5$ queues and $6$ servers. To generate such instances, we proceed as follows: generate $200$ base parameters by selecting 5 queue arrival rates, $\lambda_{i}$, and 6 server capacity rates, $\mu_{j}$, uniformly at random in $(0,1)$. For each value of the ratio $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (with $r$ in $(0,1]$), rescale the arrival rates such that the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}=r$ is satisfied. This process is repeated for each of the $200$ sets of base parameters. The simulations are run for $T=50$,$000$ and $\gamma=1/\sqrt{T}$. In Figure 2(b), the solid line represents the average buildup over the $200$ simulation for each $r$ value; the shaded region indicates the range between the $2.5$ and $97.5$ percentiles of buildup values. The results confirm that, even for asymmetric and randomized systems, the transition point remains above $0.5$ and below $1$. In both Figure 2(a) and Figure 2(b) the vertical lines indicate where $r=\frac{1}{3}$ and $r=\frac{1}{2}$. The empirical buildups in these simulations show that while the worst-case bound is $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, systems with learning agents typically do better than the worst case at clearing packets.
(a) Probability that a randomized system with time horizon $T$ has a queue with buildup greater than $\sqrt{T}$ as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$. Shaded regions show 95% confidence interval for the probability estimation.
@@
(a) Dynamics in a system with two queues with arrival rates $\lambda_{1}=1/3$, $\lambda_{2}=1/6$ and three servers with service capacities $\mu_{1}=2/3$, $\mu_{2}=2/9$, $\mu_{3}=1/9$ as a function of time.
(b) Probability of sending to server 1 in a system with 2 queues with $\lambda_{i}=1/4$, and 2 servers with $\mu_{i}=2/3$, compared to Nash equilibria. Dashed lines show time averages.
@@ 4.2 The Impact of Buffers
We now compare systems with vs. without buffers. First we note that even a buffer of just one packet at each server can significantly increase server capacity. To see this, consider a system with a single queue receiving packets at a high rate, say with probability $\lambda=\frac{1}{2}$ at each time step, and multiple servers, each with a service rate, say $\mu=\frac{1}{3}$. Without buffers, the condition for stability of such a system is that there is a single server with a service rate of $\mu>\lambda$, as a queue can only send a packet to one server at a time. However, with buffers, two such servers will make the system stable if the queue alternates sending its packets to them.
Our empirical methodology for illustrating the impact of buffers follows the approach described in the previous subsection. Figure 3(a) illustrates the effect of adding buffers to a system in the random-instance scenario, with $n=5$ queues and $m=6$ servers as described earlier. It shows the probability that some queue in the system has at time $T$ a buildup greater than $\sqrt{T}$ as a function of the arrival-to-capacity ratio for systems with and without buffers. The solid lines show the observed frequency in $300$ independent simulations for each value of $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ in $(0,1]$; the shaded region shows the $95\%$ confidence interval in estimating these probabilities. We observe that systems without buffers require a higher capacity relative to the arrival rate to prevent buildup.
Figure 3(b) shows the distributions of packet clearing rates for two identical systems, with and without buffers, over $1$,$000$ simulations with $T=10$,$000$ and $\gamma=1/\sqrt{T}$. There are $4$ queues with $\lambda_{1}=0.6,\lambda_{2}=\lambda_{3}=\lambda_{4}=0.3$ and $5$ servers with $\mu_{1}=0.8,\mu_{2}=\mu_{3}=0.4,\mu_{4}=\mu_{5}=0.2$. The clear separation between the distributions demonstrates that buffers increase efficiency. In the system without buffers, queues preferred the servers with higher capacity, which led to more competition and a lower clearing rate since the three high-capacity servers could not process all packets.444 The total service rate of the three top servers equals the total arrival rate, hence is not enough for stability. In the system with buffers, queues do learn to extract good service rates also from the lower capacity servers, which increases the overall clearing rate of the system.
@@ 4.3 Dynamics and Convergence
5 Conclusion
This work contributes to the growing literature on learning in games with carryover effects between rounds. Routing is an important setting where such effects arise and where routers use simple distributed learning algorithms. Carryover effects appear in many other strategic multi-agent settings, such as auctions with budget constraints, or in investment markets where current wealth affects future opportunities. A key observation from our results—potentially relevant more broadly to mechanism design with carryover effects—is that if the mechanism can smooth the payoffs learners experience over time, the resulting dynamics improve even with simple learners. In our model, buffers play this role by allowing a packet to be sent now and processed later. Exploring this idea in other domains is an interesting direction for future work. In the networking setting studied here, our results leave two central challenges open. The first is to close the gap between the necessary lower bound of 2 and the sufficient factor of 3 in the ratio between the total arrival and service rates required for stability. The second is to extend the analysis to more complex network topologies and general buffer sizes.